OpenStudy (anonymous):

Determine if the sequence 1, 2, 3, 5, 8, 13... is arithmetic, geometric, or niether

OpenStudy (whpalmer4):

An arithmetic sequence has a constant difference between adjacent terms. A geometric sequence has a constant ratio between adjacent terms. Check if arithmetic: 2-1 = 3-2= 5-3= 8-5= 13-8= if they are all identical, then it is arithmetic. Check if geometric: 2/1 = 3/2 = 5/3 = 8/5= 13/8= if they are all identical, then it is geometric.

hero (hero):

It's Fibonacci